OpenStudy (anonymous):

Find the hypotenuse of a right triangle whose legs are square root of 10 and square root15. =)

OpenStudy (anonymous):

\[h^2=a^2+b^2\] \[h^2=\sqrt{10}^2+\sqrt{15}^2=10+15=25\] \[h=5\]